Suppose that Apex HealthServices has four difference projects.These projects are listed below, along with the amount of capital invested and estimated corporate and market betas: Project Amount Invested Corporate beta Market Betas walk-in-clinic 500,000 1.5 1.1 MRI facility 2,000,000 1.2 1.5 Clinical laboratory 1,500,000 0.9 0.8 X-ray laboratory 1,000,000 0.5 1.0 5,000,000 (a)why do the corporate and market betas differ for the same project? (b)what is the overall corporate beta of Apex HealthServices? Is the calculated beta consistent with corporate risk theroy? (c)what is the overall market beta of Apex Health services? (d) How does the riskiness of Apex's stock compare with the riskiness of an average stock? (e) would stock investors required rate of return on Apex that is greater than, less than,or the same as the return on an average-risk stock?
Answer #1

Given:

1. Project (Walk in Clinic) - Capital = 500,000, Corporate Beta = 1.5, Market Beta = 1.1

2. Project (MRI Facility) - Capital = 20,00,000, Corporate Beta = 1.2, Market Beta = 1.5

3. Project (Clinical Laboratory) - Capital = 15,00,000, Corporate Beta = 0.9, Market Beta = 0.8

4. Project (X-Ray Laboratory) - Capital = 10,00,000, Corporate Beta = 0.5, Market Beta = 1

1) Market beta and corporate beta differs for the same project, because corporate beta tells the correlated volatility with the market and market beta tell the correlated volatility of all the projects in the market with the market.

2)Corporate beta of Apex Health Services is the weighted average of beta of all projects in the company, which is

Corporate Beta = [(5,00,000/50,00,000) X 1.5 +(20,00,000/50,00,000) X 1.2 + (15,00,000/50,00,000) X 0.9 + (10,00,000/50,00,000) X 0.5)]/1 = (0.1X1.5 + 0.4X1.2 + 0.3X0.9 + 0.2X0.5)/1 = (0.15 + 0.48 + 0.27 + 0.1)/1 =  1

Yes. it is consistent with corporate risk theory.

3) Market beta of Apex Health Services is the weighted average of beta of all projects in the company, which is

Market Beta = [(5,00,000/50,00,000) X 1.1 +(20,00,000/50,00,000) X 1.5 + (15,00,000/50,00,000) X 0.8 + (10,00,000/50,00,000) X 0.1)]/1 = (0.1X1.1 + 0.4X1.5 + 0.3X0.8 + 0.2X0.1)/1 = (0.11 + 0.6 + 0.24 + 0.02)/1 =  0.97.